RC Meeting and Run Plan, Tuesday, March 3 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection

Tuesday 3 Mar 2020.

General updates

  • Target has been running smoothly at around 53%+, but Accelerator has been having problems
    • Took ~3 hours of transverse data during Swing before beam was lost, then switched back to longitudinal
    • Only managed ~15min of longitudinal running over OWL
  • Injector did their spot move yesterday at 11am, beam back around 3pm
  • Rough beam through Swing and Owl
    • Ran OK from 3pm--6pm, then Accel down until end of Swing while experts tried to solve interception problems in the Injector [1]
      • Problems were not driven by the spot move, these are apparently the same interception issues that hit us a few days back.
    • Had a few more hours of trippy beam in OWL until beam off again at 5am
    • More work on Injector and beam back around 8am for a little production before Beam Studies program that will run through Day shift
  • Beam studies will attempt understand what is going on with the machine (beam loss, activation, etc) through Tuesday DAY shift [2]
    • Test program starts around 9am with some beam into Hall C to study Hall C beam-correlated loss in the NL and east arc (in particular)
    • Target group entered around 10am and was out at 6pm, then MCC took another couple hours to restore beam and complete their test program
    • It sounds like the Beam Studies was productive and they better understand some issues that have been plaguing the machine. They wanted to think about the data a bit and will present some options Wednesday
      • Many thanks to RadCon and OPs for sticking around a little longer while the Target folks worked to wrap up in the Hall
  • Mike Fowler rebooted the Magnet controls IOC in mid-Day shift to correct some issues
    • GUI wasn't working quite right, not sure what else may have been affected
  • Joe Did a quick walkthrough training for one person

Notes for the Future

  • 'Macro Pulse Generator' replacement in injector will halt beam operations on Wed (8:30--1:30pm)
    • Hall C may take another opportunistic access in that period (TBD)
  • Joe has two other folks that need a Hall walk through during an opportunistic down
    • Could happen Wed between 10--noon, if they respond
  • Schedule Moller measurement for Thursday DAY shift
  • 'Filler' runs for next spin up / target-down period:
    • 1 hour run with target out for measure of the background just from the beryllium windows

Pending Issues

  • pNMR issues (not sure this was looked at yet?)
  • EPR problems [3] saw some progress on Tuesday [4], but issues remain
    • Need a new fiber bundle ASAP at minimum
  • Kepco power supply commissioning/installation (Todd, Bill)
    • Bill will update status on Wed.
  • Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted [5] [6] [7] [8] [9] (Bill, Junhao)
  • Persistent BLM & ion chamber trips
    • General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
    • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
    • Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)
  • Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
    • Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device
  • OPs wished to reboot hung iochc10 [10], noted in December as well [11]
    • This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.

Opportunistic Access Jobs

Target coil rotation planned for March 13--20

  • Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program

Current Run Plan

Tuesday DAY

  • Down for beam studies for most of Tuesday DAY shift
    • Hall C will be in Controlled Access starting 10am through 2pm (likely in Laser CANS mode for a large part of it)

Tuesday SWING → Wed DAY

  • Follow the Short beam recovery procedure
    • Skip the YAG unless you are concerned about the harp scan data.
  • Take production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Wednesday DAY → Thursday OWL

  • Beam off around 08:30 for hardware work in Injector -- plan is to have beam restored by 13:30
  • When beam goes away around 8:30
    • Take longitudinal NMR
  • Wait for MCC to return beam (hopefully ~1:30pm)
  • Continue production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Thursday DAY → Friday OWL

  • Moller measurement planned for Thursday DAY shift
  • Just before handing over to Moller team
    • Take longitudinal NMR
    • Rotate target to 90deg Transverse (Target Expert only)
  • Wait for Moller to complete and get production beam re-established (late-DAY, early-SWING)
    • Follow the Short beam recovery procedure
      • Skip the YAG unless you are concerned about the harp scan data.
    • Take transverse NMR
    • Take 5 hours of production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target in transverse
    • Rotate target to 180deg Longitudinal (Target Expert only)
    • Take longitudinal NMR
  • Continue production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Standing Orders

 If beam goes away for more than 2h, or if during the downtime, MCC does operations
 that may change the beam tuning, then do [https://hallcweb.jlab.or/wiki/index.php/Beam_and_Raster_Checkout_for_A1n/d2n_(Short_Procedure)
 beam recovery procedure] (do YAG viewer part only if you are suspicious of
 the harp scan results)
  • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell the Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
  • Check the position locks are on.
  • Do your best to accommodate Hall A's request for a PITA scan promptly. It is OK to stop the present runs and start new ones in 5--10 minutes after they are done.
  • Check that CODA is running normally (had one strange hangup lately)
  • Check online replays shortly after starting runs, & full replay results too.
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • Ramp rate lock instructions are here and added to checklist[12]

Bluejeans remote connection information

 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Target Polarization Angle Conventions

Be VERY clear here about the convention (beam-right/left/up/downstream) vs 0, 90, 180, 270)
for each run period.  More notes and references to log entries are better.

Polarization Calibration Data

ADD LIST OF CALIBRATION RUNS HERE
AND A TABLE OF RELEVANT CALIBRATION CONSTANTS TO USE FOR WHAT RUNS
Cell Calibration
Type
SHMS Angle
/degree
SHMS Momentum
/GeV
Field Direction
/degree
Calibration Results Status Oven Temperature
/°C
Log#
Tommy EPR-NMR 11 7.5 90 5.24 %/mV Done 200 3839939
Tommy EPR-NMR 18 5.6 270 5.04(1) %/mV Done 205 3843301
Tommy EPR-NMR 18 5.6 90 undone 205
Tommy EPR-NMR 11 7.5 270 undone 208
Tommy AFP Loss 14.5 6.4 90 1.25%/Sweep Done 200 3840618, 3840143, 3840618
Tommy EPR Amplifier
on/off NMR Signal
11 7.5 90 Done 200 3839962
Tommy Temperature Test 18 5.6 90 Done 205 3844089
Tommy AFP Loss 18 5.6 270 Done 205 3845255
Tommy Temperature Test 18 5.6 270 Done 205 3845336
Briana EPR-NMR 18 5.6 180 8.60 Done 205 3834036, 3834243
Briana EPR-NMR 18 5.6 270 5.6 Done 205 3835512, 3835587
Austin EPR-NMR 18 5.6 90 4.88 Done 200 3814717
Austin EPR-NMR 11 7.5 180 7.64 Done 210 3817200
Austin EPR-NMR 11 7.5 0 8.13 Done 210 3818761
Austin EPR-NMR 11 7.5 270 4.56 Done 210 3819882
Austin EPR-NMR 11 7.5 270 4.58 Done 210 3822296
Austin EPR-NMR 11 7.5 90 4.91 Done 210 3824907
Austin EPR-NMR 18 0 90 4.58 Done 210 3815895
Austin EPR-NMR 11 0 180 6.91 Done 210 3816724

Target Polarization Database

d2n Run Period

D2n Target Run List: https://docs.google.com/spreadsheets/d/16TrIUXVA-znVNUyEKy9FHaIlcvq-KXzu97k9nJowuE8/edit#gid=0

  • Junhao Chen is maintaining it

A1n Run Period

https://docs.google.com/spreadsheets/d/e/2PACX-1vSoTyDugD8wt3hKmBGZBxwUJTPHmUgyvGVJvM7N6qeVoNWhEXumCwwkPI7Sw9Asae3N4k9h5L0Y7qmw/pubhtml

  • Mingyu Chen will updating the runlist while Murchhana Roy will do the cross check.

Recent polarization measurements

PolHe3 Polarization Measurements






Misc Links